# Functional differential equations driven by càdlàg rough paths

### v2

@@ Appendix A Local estimates for rough integration
For $p\in(2,3)$, suppose $\mathbf{X},\tilde{\mathbf{X}}\in\mathcal{D}^{p}([0,T];\mathbb{R}^{d})$, $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$, $(\tilde{Y},\tilde{Y}^{\prime})~{}\in\mathcal{V}^{p}_{\tilde{X}}([0,T];\mathbb{R}^{k})$, and that the non-anticipative functional $(F,F^{\prime})\colon\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})\to\mathcal{V}^{p}_{X}([0,T];\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$ satisfies Assumption 2.1. Then, we have the local estimate
$\|R^{\int_{0}^{\cdot}F(Y)\mathrm{d}\mathbf{X}}-R^{\int_{0}^{\cdot}F(\tilde{Y})\mathrm{d}\tilde{\mathbf{X}}}\|_{\frac{p}{2},[s,t]}$
$\quad\lesssim C_{F,K,X,\tilde{X}}(|Y_{s}-\tilde{Y}_{s}|+\|Y,Y^{\prime};\tilde{Y},\tilde{Y}^{\prime}\|_{X,\tilde{X},p,[s,t]}+\|X-\tilde{X}\|_{p,[s,t]})(\|\mathbf{X}\|_{p,[s,t]}\vee\|\tilde{\mathbf{X}}\|_{p,[s,t]})$
$\qquad+C_{F}(1+K)^{2}(1+\|X\|_{p,[s,t]}\vee\|\tilde{X}\|_{p,[s,t]})^{2}\|\mathbf{X};\tilde{\mathbf{X}}\|_{p,[s,t]}$
for all $(s,t)\in\Delta_{T}$, if $\|Y,Y^{\prime}\|_{X,p,[s,t]}$, $\|\tilde{Y},\tilde{Y}^{\prime}\|_{\tilde{X},p,[s,t]}\leq K$, for some $K>0$, where the implicit multiplicative constant depends on $p$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$.
@@
It follows from [FZ18, Lemma 3.4] that for any $(V,V^{\prime})\in\mathcal{V}^{p}_{X}$, $(\tilde{V},\tilde{V}^{\prime})\in\mathcal{V}^{p}_{\tilde{X}}$,
$\|R^{\int_{0}^{\cdot}V\mathrm{d}\mathbf{X}}-R^{\int_{0}^{\cdot}\tilde{V}\mathrm{d}\tilde{\mathbf{X}}}\|_{\frac{p}{2},[s,t]}$
$\quad\lesssim_{p}(1+\|\mathbf{X}\|_{p,[s,t]}+\|\tilde{\mathbf{X}}\|_{p,[s,t]})(\|V,V^{\prime};\tilde{V},\tilde{V}^{\prime}\|_{X,\tilde{X},p,[s,t]}\|\mathbf{X}\|_{p,[s,t]}$
$\qquad+\|\tilde{V},\tilde{V}^{\prime}\|_{\tilde{X},p,[s,t]}\|\mathbf{X};\tilde{\mathbf{X}}\|_{p,[s,t]})$
$\quad\lesssim\|V,V^{\prime};\tilde{V},\tilde{V}^{\prime}\|_{X,\tilde{X},p,[s,t]}\|\mathbf{X}\|_{p,[s,t]}+\|\tilde{V},\tilde{V}^{\prime}\|_{\tilde{X},p,[s,t]}\|\mathbf{X};\tilde{\mathbf{X}}\|_{p,[s,t]},$
where the implicit multiplicative constant depends on $p$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$. For $(V,V^{\prime})=(F(Y),F^{\prime}(Y,Y^{\prime}))$, $(\tilde{V},\tilde{V}^{\prime})=(F(\tilde{Y}),F^{\prime}(\tilde{Y},\tilde{Y}^{\prime}))$, using Assumption 2.1 (ii), we therefore obtain the estimate. ∎

### v3

@@ Appendix A Local estimates for rough integration
For $p\in(2,3)$, suppose $\mathbf{X},\widetilde{\mathbf{X}}\in\mathcal{D}^{p}([0,T];\mathbb{R}^{d})$, $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$, $(\widetilde{Y},\widetilde{Y}^{\prime})~{}\in\mathcal{V}^{p}_{\widetilde{X}}([0,T];\mathbb{R}^{k})$, and that the non-anticipative functional $(F,F^{\prime})\colon\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})\to\mathcal{V}^{p}_{X}([0,T];\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$ satisfies Assumption 2.4 (i) and (ii) given $X,\widetilde{X}$. Then, we have the local estimate
$\|R^{\int_{0}^{\cdot}F(Y)\mathrm{d}\mathbf{X}}-R^{\int_{0}^{\cdot}F(\widetilde{Y})\mathrm{d}\widetilde{\mathbf{X}}}\|_{\frac{p}{2},[s,t]}$
$\quad\lesssim C_{F,K,X,\widetilde{X}}(|Y_{s}-\widetilde{Y}_{s}|+\|Y,Y^{\prime};\widetilde{Y},\widetilde{Y}^{\prime}\|_{X,\widetilde{X},p,[s,t]}+\|X-\widetilde{X}\|_{p,[s,t]})(\|\mathbf{X}\|_{p,[s,t]}\vee\|\widetilde{\mathbf{X}}\|_{p,[s,t]})$
$\qquad+C_{F}(1+K)^{2}(1+\|X\|_{p,[s,t]}\vee\|\widetilde{X}\|_{p,[s,t]})^{2}\|\mathbf{X};\widetilde{\mathbf{X}}\|_{p,[s,t]}$
for all $(s,t)\in\Delta_{T}$, if $\|Y,Y^{\prime}\|_{X,p,[s,t]}$, $\|\widetilde{Y},\widetilde{Y}^{\prime}\|_{\widetilde{X},p,[s,t]}\leq K$, for some $K>0$, where the implicit multiplicative constant depends on $p$, $\|\mathbf{X}\|_{p}$ and $\|\widetilde{\mathbf{X}}\|_{p}$.
@@
It follows from [FZ18, Lemma 3.4] that for any $(V,V^{\prime})\in\mathcal{V}^{p}_{X}$, $(\widetilde{V},\widetilde{V}^{\prime})\in\mathcal{V}^{p}_{\widetilde{X}}$,
$\|R^{\int_{0}^{\cdot}V\mathrm{d}\mathbf{X}}-R^{\int_{0}^{\cdot}\widetilde{V}\mathrm{d}\widetilde{\mathbf{X}}}\|_{\frac{p}{2},[s,t]}$
$\quad\lesssim_{p}(1+\|\mathbf{X}\|_{p,[s,t]}+\|\widetilde{\mathbf{X}}\|_{p,[s,t]})(\|V,V^{\prime};\widetilde{V},\widetilde{V}^{\prime}\|_{X,\widetilde{X},p,[s,t]}\|\mathbf{X}\|_{p,[s,t]}$
$\qquad+\|\widetilde{V},\widetilde{V}^{\prime}\|_{\widetilde{X},p,[s,t]}\|\mathbf{X};\widetilde{\mathbf{X}}\|_{p,[s,t]})$
$\quad\lesssim\|V,V^{\prime};\widetilde{V},\widetilde{V}^{\prime}\|_{X,\widetilde{X},p,[s,t]}\|\mathbf{X}\|_{p,[s,t]}+\|\widetilde{V},\widetilde{V}^{\prime}\|_{\widetilde{X},p,[s,t]}\|\mathbf{X};\widetilde{\mathbf{X}}\|_{p,[s,t]},$
where the implicit multiplicative constant depends on $p$, $\|\mathbf{X}\|_{p}$ and $\|\widetilde{\mathbf{X}}\|_{p}$. For $(V,V^{\prime})=(F(Y),F^{\prime}(Y,Y^{\prime}))$, $(\widetilde{V},\widetilde{V}^{\prime})=(F(\widetilde{Y}),F^{\prime}(\widetilde{Y},\widetilde{Y}^{\prime}))$, using Assumption 2.4 (ii), we therefore obtain the estimate. ∎